Solve the stem of linear equations. What is the x and y answer to this problem?
y - 2x = 0
6y + 2x = 0

Answer by mananth(16946)

-2 x + 1 y = 0
2 x + 6 y = 0
Eliminate y
multiply (1)by -6
Multiply (2) by 1
12 x -6 y = 0
2 x 6 y = 0
Add the two equations
14 x = 0
/ 14
x = 0
plug value of x in (1)
-2 x + 1 y = 0
0 + 1 y = 0
1 y = 0
1 y = 0
y = 0